OpenStudy (anonymous):

use additive inverse to solve 3x+25=56 what do you do first

OpenStudy (anonymous):

Subtract a constant from both sides to start getting x by itself. Or another way to put it is add the opposite of the constant.

OpenStudy (anonymous):

3x+25=56 -25 -25 3x=31 x=31/3

OpenStudy (anonymous):

my answer choices are a- add 25 to both sides b- add -25 to both sides c- add 3 to both sides d-add -3 to both sides

OpenStudy (anonymous):

so what would my answer be

OpenStudy (anonymous):

B

OpenStudy (anonymous):

What's the opposite of the constant?

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

Do you know what a constant is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a constant is a number on its own

OpenStudy (anonymous):

OK, then we need to start there. A constant is a number that is what it is. There's nothing changing it. It's constantly the same. A variable is a number that is having something done to it. Like 3x. 3x isn't a 3. It's a 3 times something else. What it really is depends on what that x is. If x is 2, then it's a 6. If x is 3, then it's a 9. It varies, hence variable.

OpenStudy (anonymous):

ok so how is the answer b if 56 - 25 is 31 and 56 + 25 is 81

OpenStudy (anonymous):

so you -25 to the 56 because your trying to get the x by itself so your subtracting the constant that's on the same side of the variable. the answer is b.

OpenStudy (anonymous):

Because of what the question is asking you. It's asking you what you do FIRST, not what the result of doing it is.

OpenStudy (anonymous):

ok so you add negative 25 first

OpenStudy (anonymous):

Exactly. Because your goal is to get anything with x in it on one side and the constants on the other side.

OpenStudy (anonymous):

yes to the 53

OpenStudy (anonymous):

ok thanks can yall help me out somemore please witt a couple other problems

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how do i solve this inequality \[15 <5 - 2 \left( 4m \+7)\]

OpenStudy (anonymous):

15< 5 - 2 (4m+7)

OpenStudy (anonymous):